OpenStudy (anonymous):

what is the solving equation for "5x+3 =23"???

OpenStudy (anonymous):

5x+3=23 23=5x+3 23-3=5x 20=5x x=20/5 x=1

OpenStudy (anonymous):

are you trying to find x?

OpenStudy (anonymous):

sorry x=4